# LCDB 1.1: A Database Illustrating Learning Curves Are More Ill-Behaved Than Previously Thought

### v1

@@ E.3 More Details regarding LDA
Mathematical Proof
@@
Proof: The scaling factor for feature scaling transformation $\mathbf{S}$ is a diagonal matrix, after feature scaling $\mathbf{x}^{\prime}=\mathbf{S}\mathbf{x}$. Correspondingly, the class means and covariance matrices become $\bm{\mu}_{k}^{\prime}=\mathbf{S}\bm{\mu}_{k}$ and $\bm{\Sigma}^{\prime}=\mathbf{S}\bm{\Sigma}\mathbf{S}$.
@@
Step 18 since $\mathbf{S}$ is diagonal. The new discriminant function is the same as no feature scaling one, so the learning curves should be exactly the same after feature scaling.
Note that a similar argument holds for QDA, which should also be insensitive to feature scaling. However, given the reproducibility issues of QDA we did not investigate this.
E.4 Detailed Violations Error in CC-18
Since we define the violation error in Equations 2 and 3 to quantify the size of monotonicity and convexity violations, we are able to compare the severity of such violations across different learners and datasets. Figures 13 and 14 provide heatmap visualizations illustrating these violation errors, where zero means no violation, and white means missing learning curves.
@@ Appendix F Learning Curves Fitting with More Parametric Models
[…]
@@
Figure 15: Ill-behaved (non-monotone or non-convex) learning curves are more difficult to fit with parametric models. MSE of the fitting is proportional to the violation error. (a, b) Models: MMF4 ($\frac{ab+cx^{d}}{b+x^{d}}$). (c, d) WBL4 ($-b\exp(-ax^{d})+c$).
[…]
@@ Appendix G Learning Curve Crossings Affect Successive Halving
To showcase how we can use LCDB 1.1 to study the connection between SH model selection performance and crossing of learning curves, we show more detailed experimental results for comparing the performance of SH on two groups of learners, one with rarely and one with commonly crossing learning curves. In Figure  18, the blue and orange groups are the learners whose curves cross rarely and frequently, respectively. The left, middle, and right column figures are for the case where SH starts at the first available anchor (16 training instances), the 8th anchor (30 training instances), and the 16th anchor (59 training instances), respectively. From top to bottom, each row corresponds to a different per-round budget increase rate in the SH procedure, specifically 12.5%, 25.0%, 50.0%, and 100%. The per-round budget increase rate determines how much the training budget is increased between consecutive SH rounds. The left panel of box-plots show, for different values of $k$, the probability (across the 5 outer folds of datasets) that the finally chosen algorithm is under the top $k$. However, since the final performance differences of learners may be small, we complement this figure with the regrets on the right panel of box-plots, i.e., final error rate of the chosen learner minus the minimum of final error rate (in the log-scale).
@@ Appendix H Alternative Method To Detect Monotonicity Violations
Appendix I Broader Discussion about LCDB 1.1
@@
The creation of LCDB 1.1 involved approximately 600,000 CPU-hours of computing. The computing was conducted on a heterogeneous cluster environment, with jobs scheduled across a range of CPU nodes. Since the specific CPU node used for each job was not fixed, we provide a representative list of the hardware types used during experimentation: AMD EPYC 7502P 32-Core Processor, AMD EPYC 9534 64-Core Processor, AMD EPYC 7413 24-Core Processor, AMD EPYC 7452 32-Core Processor, AMD EPYC 7543 32-Core Processor, Intel(R) Xeon(R) CPU E5-2667 v4 @ 3.20GHz, Intel(R) Xeon(R) CPU E5-2680 v4 @ 2.40GHz, Intel(R) Xeon(R) Gold 5218 CPU @ 2.30GHz, Intel(R) Xeon(R) Gold 6130 CPU @ 2.10GHz, Intel(R) Xeon(R) Gold 6140 CPU @ 2.30GHz. Each compute job was allocated 20 GB of memory. We set the maximum execution time as 3.5 hours for computing per anchor, inner, and outer seed combination. If the total runtime exceeds this limit, the job is terminated and its status is recorded as a timeout. Both the error message and execution status were logged.

### v2

@@ E.4 More Details regarding LDA
To theoretically confirm that feature scaling does not affect the LDA learning curves, we provide the following proof.
Proof.
@@
$\mathbf{x}$ is the feature vector, $\bm{\mu}_{k}$ and $\bm{\Sigma}$ denote the class mean and shared covariance matrix, and $p(y_{k})$ is the prior probability of class $k$.
The scaling factor for feature scaling transformation $\mathbf{S}$ is a diagonal matrix, after feature scaling $\mathbf{x}^{\prime}=\mathbf{S}\mathbf{x}$. Correspondingly, the class means and covariance matrices become $\bm{\mu}_{k}^{\prime}=\mathbf{S}\bm{\mu}_{k}$ and $\bm{\Sigma}^{\prime}=\mathbf{S}\bm{\Sigma}\mathbf{S}$.
@@
The new discriminant function is the same as no feature scaling one, so the learning curves should be exactly the same after feature scaling. ∎
Note that a similar argument holds for QDA, which should also be insensitive to feature scaling. However, given the reproducibility issues of QDA we did not investigate empirically.
E.5 Detailed Violations Error in CC-18
Since we define the violation error in Equations 2 and 3 to quantify the size of monotonicity and convexity violations, we are able to compare the severity of such violations across different learners and datasets. Figures 15 and 16 provide heatmap visualizations illustrating these violation errors, where zero means no violation, and white means missing learning curves.
@@ Appendix F Learning Curves Fitting with More Parametric Models
To further validate the findings in Section 5.4, we include additional experiments with parametric models $\mathrm{MMF4}$ and $\mathrm{WBL4}$ by using the same experimental setting (Figure 17). These results confirm that the conclusions remain consistent.
@@
Figure 17: Ill-behaved (non-monotone or non-convex) learning curves are more difficult to fit with parametric models. MSE of the fitting is proportional to the violation error. (a, b) Models: MMF4 ($\frac{ab+cx^{d}}{b+x^{d}}$). (c, d) WBL4 ($b\exp(-ax^{d})+c$).
The intrinsic properties of these parametric models conflict with the characteristics of some ill-behaved learning curves, which explains the observed experimental results. In particular, the phase transition shapes (only the convexity violated) can be effectively fitted by $\mathrm{MMF4}$. The peaking and dipping, which violate the monotonicity, cannot be modeled by these parametric models. For reference, we also provide illustrative examples (Figure 18) highlighting the alignment (or misalignment) between model properties and the geometric characteristics of learning curves, as discussed in the main text.
@@ Appendix G Learning Curve Crossings Affect Successive Halving
To showcase how we can use LCDB 1.1 to study the connection between SH model selection performance and crossing of learning curves, we show more detailed experimental results for comparing the performance of SH on two groups of learners, one with rarely and one with commonly crossing learning curves. In Figure  20, the blue and orange groups are the learners whose curves cross rarely and frequently, respectively. The left, middle, and right column figures are for the case where SH starts at the first available anchor (16 training instances), the 8th anchor (30 training instances), and the 16th anchor (59 training instances), respectively. From top to bottom, each row corresponds to a different per-round budget increase rate in the SH procedure, specifically 12.5%, 25.0%, 50.0%, and 100%. The per-round budget increase rate determines how much the training budget is increased between consecutive SH rounds. The left panel of box-plots show, for different values of $k$, the probability (across the 5 outer folds of datasets) that the finally chosen algorithm is under the top $k$. However, since the final performance differences of learners may be small, we complement this figure with the regrets on the right panel of box-plots, i.e., final error rate of the chosen learner minus the minimum of final error rate (in the log-scale).
@@ Appendix H Alternative Method To Detect Monotonicity Violations
Appendix I Statistical Correction Using Holm’s Method
We consider to use a slightly less conservative method called Holm’s Step-Down Procedure (Holm’s method) [18], in the sense that it will reject more null hypotheses, typically resulting in fewer Type II errors. As shown in Table 9, the results remain consistent with our original findings: an even larger fraction of learning curves are identified as ill-behaved, up to 19%.
Table 9: Ill-behavior statistics of the LCDB 1.1 (23 learners, Dummy excluded) by using Holm’s method.
Shapes / Database LCDB 1.1 CC-18 (72) LCDB 1.1 FULL (265)
no FS min-max FS standardization FS no FS min-max FS standardization FS
Missing 1.9% 0.0% 8.7% 2.6% 0.0% 8.7%
Non-Monotone ($\neg$ M) 13.3% 13.0% 10.6% 13.2% 14.1% 11.8%
Non-Convex ($\neg$ C) 14.6% 14.1% 11.8% 14.8% 15.4% 13.5%
Ill-behaved ($\neg$ M $\cup$ $\neg$ C) 18.2% 17.3% 14.7% 18.8% 19.8% 16.8%
Appendix J Broader Discussion about LCDB 1.1
@@ Appendix I Statistical Correction Using Holm’s Method
The creation of LCDB 1.1 involved approximately 800,000 CPU-hours and 3,000 GPU-hours of computing. The computing was conducted on a heterogeneous cluster environment provided by the Delft Artificial Intelligence Cluster (DAIC) [77]. Since the specific CPU node used for each job was not fixed, jobs were scheduled across a range of CPU nodes: AMD EPYC 7502P 32-Core Processor, AMD EPYC 9534 64-Core Processor, AMD EPYC 7413 24-Core Processor, AMD EPYC 7452 32-Core Processor, AMD EPYC 7543 32-Core Processor, Intel(R) Xeon(R) CPU E5-2667 v4 @ 3.20GHz, Intel(R) Xeon(R) CPU E5-2680 v4 @ 2.40GHz, Intel(R) Xeon(R) Gold 5218 CPU @ 2.30GHz, Intel(R) Xeon(R) Gold 6130 CPU @ 2.10GHz, Intel(R) Xeon(R) Gold 6140 CPU @ 2.30GHz. For TabPFN v2 learning curves, we used DelftBlue Supercomputer hosted at the Delft High Performance Computing Centre (DHPC) [78], utilizing Nvidia A100 and Nvidia Tesla V100. Each compute job was allocated 20 GB of memory. We set the maximum execution time as 3.5 hours for computing per anchor, inner, and outer seed combination. If the total runtime exceeds this limit, the job is terminated and its status is recorded as a timeout. Both the error message and execution status were logged.